OpenStudy (anonymous):

What is the exact value of the expression:

OpenStudy (anonymous):

\[7\sqrt{8} + 2\sqrt{12} - 3\sqrt{2}\]

OpenStudy (anonymous):

@FutureMathProfessor

OpenStudy (anonymous):

Break down the 8 and the 12 inside the square roots :)

OpenStudy (anonymous):

make them 2?

OpenStudy (anonymous):

@FutureMathProfessor

OpenStudy (anonymous):

\[7\sqrt{8} + 2\sqrt{12} - 3\sqrt{2}\] \[=7\sqrt{4\times 2}+2\sqrt{4\times 3}-3\sqrt{2}\] \[=7\sqrt4\sqrt2+\sqrt4\sqrt3-3\sqrt2\] \[=7\times 2\sqrt2+2\times 2\sqrt3-3\sqrt2\] is a start then clean it up some with arithmetic and combining like terms

OpenStudy (anonymous):

\[2\sqrt{4}\sqrt{3}\] is it supposed to be that for the second one? @satellite73